Y=2x+1
X+y=7 this is a substitution i gotta find the x and y

Answer by rothauserc(4718)

y = 2x+1
y = 7-x
substitute for y in first equation
7-x = 2x+1
3x = 6
x = 2
y = 7-2 = 5
therefore
x = 2 and y = 5